OpenStudy (anonymous):

there are 1.806x10^25 molecules of water (H20). how many moles of water does the sample contain? a: 30.0 moles b: 90.0 moles c: 1.08 x 10^48 moles d: 6.02 x 10^24 moles

sam (.sam.):

\[1 mol= 6.02x10^{23}atomss\] There are 3 atoms in one molecule, \[1.806molecules \times\frac{3atoms}{1molecule}=5.418 \times10^{25}atoms\] \[5.418 \times10^{25}atoms \times\frac{1mole}{6.02 \times 10^{23}atoms}=90moles\]